OpenStudy (anonymous):

A 10.0 kg block is sliding at 8.94 m/s in a straight line across a horizontal, frictionless surface. The block then passes over a 4.00 m stretch of surface with friction. The kinetic coefficient of friction of this patch is .600. What is the velocity of the block after it passes over the section of surface with friction?

OpenStudy (shamim):

Frictional force=coefficient of friction*m*g=.6*10*9.8=?

OpenStudy (shamim):

Now u hv to use the formula F=ma .6*10*9.8=10*a a=?

OpenStudy (shamim):

Now u hv to use v=u+at v=?

OpenStudy (shamim):

Feel free to ask for more clarification

OpenStudy (anonymous):

what do you plug in for u and t?

OpenStudy (shamim):

Sorry, u hv to use v^2=u^2-2*a*s Put u=8.94m/s s=4m a=.6*9.8m/s^2

OpenStudy (shamim):

v=?

OpenStudy (anonymous):

got it thanks a bunch!!

OpenStudy (shamim):

Glad to help u!!!!